OpenStudy (dimpalk104):

~~will medal and fan~~~ At a particular pressure and temperature, it takes just 5.612 min for a 1.057 L sample of Cl2 to effuse through a porous membrane. How long would it take for the same volume of UF4 to effuse under the same conditions? ???min

OpenStudy (aaronq):

You have to use graham's law of effusion: \(\sf \dfrac{rate_A}{rate_B}=\sqrt{\dfrac{M_B}{M_A}}\) where M is the molar mass of a substance. To find the time, we use the rate. \(\sf rate=\dfrac{molecules}{time }\) substitute this back into the equation: \(\sf \dfrac{\dfrac{molecules_A}{time_A }}{\dfrac{molecules_B}{time_B }}=\sqrt{\dfrac{M_B}{M_A}}\) Were told the same volume (i.e. number of molecules) effuses through, so: \(\sf molecules_A=molecules_B\) The equation reduces to: \(\sf \dfrac{time_B}{time_A}=\sqrt{\dfrac{M_B}{M_A}}\) Solve for time of UF_4: \(\sf time_{UF_4}=time_{Cl_2}*\sqrt{\dfrac{M_{UF_4}}{M_{Cl_2}}}\)